uneven cash flow stream:

What are the PVs of the streams at 0% compounded annually?


0-------------1----------------2---------------3----------------4-------------5

stream a $0 $100 $400 $400 $400 $300

stream b $0 $300 $400 $400 $400 $100


1
Expert's answer
2021-05-10T12:53:46-0400

"Stream A"

"Present { }value = (100\/1.00^1)+(400\/1.00^2)+"

"(400\/1.00^3)+(400{\/}1.00^4)+(300\/1.00^5)"

"100+400+400+400+300 = 1600Dollars""Presentvalue B=(300\/1.00^1 )+(400\/1.00^2)+"

"(400\/1.00^3)+(400\/1.00^4)+(100\/1.00^5)="

"300+400+400+400+100= 1600 Dollars"